OpenStudy (anonymous):

help: Find the slope and the y-intercept of the equation 2x − 3y = 6. slope = 2/3 and y-intercept = −2 slope = −2/3 and y-intercept = −2 slope = 2/3 and y-intercept = 2 slope = −2/3 and y-intercept = 2

OpenStudy (anonymous):

Move the 2x to other side by subtracting -2x from both sides. -3y = -2x + 6 Divide through by -3 y = (2/3)x - 2 m = slope = 2/3 b = y intercept = -2, or the point (0, -2)

OpenStudy (anonymous):

2x + 3y = 1,470 3y = -2x + 1470 y = (-2x + 1470 ) / 3 y = -2/3 * x + 1470/3 y = -2/3 * x + 490 slope is -2/3 y intercept is 490 TRY USING THIS METHOD JUST NOT THE SAME NUMBER OBVIOUSLY

OpenStudy (anonymous):

guys remember not to give direct answers @LavinaKalra and @Young_gliss

OpenStudy (anonymous):

re-arrange the terms to change it to a form like y = mx +c where m is slope and c is y intercept this is done as: 2x-3y =6 so, just rearranging the equation a bit => -3y = 6-2x or 3y = 2x-6 divide by 3 so this equation will turn into => y = (2/3)x -2 so clearly slope is 2/3 and intercept is -2